# Modular Jump Gaussian Processes

### v2

@@ 4.2 A new feature for GP regression
This new variable should be related to component membership. We propose using the probability of component membership [Section 4.1], which should be high for observations likely to be in one component and low for the other. E.g., classifier $\hat{f}(x)$ predicts the probability each location $x$ belongs to the red region. I.e., it is large (close to 1) in deeply red regions and small (close to 0) in deeply yellow regions. Suppose we treat those values as new feature, comprising a new column of $X_{N}$ for GP modeling. In the new, higher dimensional input space, points in the red region will be closer together than those from the yellow region, and vice versa. Under the likelihood (2) and when predicting (4), points from the same component will have more influence on each other than on points from opposite components.
@@
To illustrate the effect this new feature has, consider the center plot of Figure 7, again focusing on x=(0.32,0.17)x=(-0.32,0.17) and its nearest neighbor $x_{*}$. We also plot the 160 NNs to $x$, color coded by the rank of their distance from $x$. The points that are the lightest are closest to $x$, and those that are the darkest are furthest away. Note that these distances are calculated prior to hyperparameter fitting. Observe first that the neighborhood is no longer circular; it has become oval-shaped to avoid the yellow region. Any numbered points that are in the yellow region are much further away than their 2d distance would suggest. Their colors are darker than nearby points the red region, and thus farther from $x$.
@@
The right panel of Figure 7 provides a MJGP fit on the Phantom data, contrasting with Figure 4. In this case the GP used is an OLAGP, chosen to illustrate the change in neighborhood structure induced by the MJGP. Here, in Figure 7, both neighborhoods are smaller, but are completely contained to the same side of the manifold. MJGP has visually improved prediction both visually and via RMSPE (see Appendix B.1) over the global GP, LAGP, and OLAGP.
Now return to Figure 5, summarizing all 10K-odd $\hat{n}(x)$ for OLAGP and MJGP. Again, the choice of $n=50$ (the LAGP default) is not particularly advantageous, except for a small percentage of $x\in\mathcal{X}$. Notice that MJGP slightly prefers more mid-range neighborhood sizes than the optimal neighbors GP. In particular, many fewer small $\hat{n}(x)$ are chosen since the neighborhood shape is no longer circular in the original input space (i.e., ignoring the new feature). This facilitates excluding points from the opposite component without resorting to extremely small neighborhood sizes.
@@ 5 Implementation and empirical benchmarking
All experiments were completed in R. Code is provided at https://bitbucket.org/gramacylab/jumpgp. In this code we provide functions aGP.optneigh.par(), mjgp(), and jgp() to run OLAGP, MJGP, and OJGP, respectively. OJGP leveraged R.matlab (Bengtsson,, 2022).
@@
OLAGP is parallelized over predictive locations $x\in\mathcal{X}$, turning statistical independence into computational exclusivity via foreach (Microsoft and Weston,, 2022). For MJGP, we utilize EM-based clustering and either a CGP (Williams and Barber,, 1998) or a random forest (RF; Breiman,, 2001) for classification via mclust (Scrucca et al.,, 2023), kernlab (Karatzoglou et al.,, 2004) and randomForest (Liaw and Wiener,, 2002), respectively. We only prefer a RF over a CGP when $n$ is large, for example during the Modified Michalewicz experiment. Again, see Table 1. Model performance is measured out-of-sample using $\mathrm{RMSPE}=(\frac{1}{m}\sum_{x\in\mathcal{X}}(y(x)-\mu(x))^{2})^{1/2},$ for which lower is better. Computation times are in Appendix D and prediction interval coverage is in Appendix E.
@@ 5.1 Phantom
For a more granular view, the right panel draws black lines connecting RMSPEs from the same rep, allowing us to see the best performing method for every training/testing set combination. We conclude that most of the overlap of the boxplots is due to variability in the training/testing split. The ordering of accuracy from one MC rep to the next is consistent across methods.
In these plots (and all plots in this section), our two methods are shown third from right (OLAGP) and furthest right (MJGP). Both show improvement over the global GP and LAGP, although OLAGP is typically outperformed by OJGP. This is unsurprising, since OJGP was designed specifically for data from jump processes, and OLAGP was not. But notice that MJGP consistently outperforms OJGP, here and broadly over the next few subsections. Looking first at the line plot on the right of Figure 8, we see that for every data set split, predictions made by MJGP are more accurate than those from OJGP. Additionally, the boxplot shows that OJGP has higher variability than MJGP. This added variability is explained by the use of a quadratic rather than linear boundary; for more on this distinction, see Park, (2022).
@@ 5.2 Star
Now consider the “Star” data, which we created for this paper. The goal was to have something like the Phantom, but more angular. Details for how the data set was created are provided in Appendix A.2. See the left panel of Figure 9, with marginal response histogram depicted earlier in the left panel of Figure 6. Like the Phantom, the Star has two well-separated regimes that form a clear shape on the response surface. But the Star is more difficult to model because all points in the yellow regime are geographically close to orange ones. It is hard to build a neighborhood inside the yellow portion of the star.
@@
Repeating themes from the Phantom analysis, consider model performance for a single Star train–test split. Mean fits for the global GP and LAGP are shown in Figure 17 and in Appendix B.2, whereas those of OLAGP and MJGP are shown here in the right two panels of Figure 9. RMSPEs can be found in Appendix B.2, where we see that the local methods perform well, with MJGP producing the best predictions of the four. The scenarios depicted in right two panels focus on $x=(0.15,0.8)$ and $(0.83,0.11)$. Compared to LAGP, OLAGP has chosen a neighborhood of roughly the same size for $(0.83,0.11)$, and much smaller for $(0.15,0.8)$; however, this smaller neighborhood still contains points from the orange regime. MJGP provides an interesting change: a much smaller neighborhood size for $(0.15,0.11)$, and a larger neighborhood size for $(0.83,0.11)$. Both neighborhoods have adapted their shape to only include points from the same regime, and this improves prediction overall.
@@ 5.3 Modified Michalewicz
Since $N$ here is very large, we must omit the global GP as a competitor in this experiment. Instead of a CGP we deploy a RF as our classifier. Finally, we limit the number of MC reps to thirty. Results are provided in Figure 12. LAGP, OLAGP, and MJGP all considerably outperform OJGP. LAGP and OLAGP perform comparably, although interestingly LAGP seems to slightly outperform OLAGP in this case. We suspect this is due to the spiky nature of the response surface; see Figure 11. MJGP performs best of all, yielding the lowest RMSPE in 29 of the 30 reps.
@@ 5.4 AMHV transport
Finally, consider the AMHV transport data first introduced in Section 1. Specifically, we used the larger version of the data set described in Section 5.2 of Park et al., (2023), scaled so that the inputs lie in the unit hypercube; $d=4$ and N=9,503N=9{,}503. A visual is provided in the right panel of Figure 11. Although $N$ is similar in magnitude to the sample size of the Star and Phantom data sets (N=10,201)(N=10{,}201), the addition of two dimensions means that there is a larger volume to fill. In that sense, this example is the most sparse of the ones entertained in this paper. Outputs are derived from a simulation of the behavior of AMHVs in a particular facility, tallying the average transport time of the AMHVs, including any time spent waiting for the vehicle to arrive. Park et al., (2023) show that as production in the facility increases, transport time, which depends on various design variables, increases and becomes a major bottleneck to the entire manufacturing process. The four variables here include vehicle speed, vehicle acceleration, minimum distance required between vehicles, and the maximum search range of empty vehicles.
@@
The results of the experiment are shown in Figure 13. As expected, the local methods do not perform as well as the global ones, with OLAGP suffering most. But the global GP performs well, producing comparable results to OJGP. MJGP produces only slightly better results, but is consistent in doing so (lowest MSPEs in 99 of the 100 reps).
@@ B.1 Phantom
Both fits are visually less accurate than the OLAGP fit in Figure 4 and MJGP fit in Figure 7. This is verified by the table of RMSPE values in the right panel, which shows that both OLAGP and MJGP produce more accurate predictions than either method. The LAGP fit in the center panel also shows the selected neighborhoods for the same two predictive locations used in Sections 3 and 4, (0.32,0.17)(-0.32,0.17) and $(0.39,0.34)$. Notice in particular that the neighborhood for $(0.39,0.34)$ contains points in both the orange and yellow regions, which explains why our prediction is inaccurate. We are able to select smaller and differently shaped neighborhoods using OLAGP and MJGP so that our predictions are more accurate. We can verify that OLAGP and MJGP make more accurate predictions by looking at Figure 16, which shows the absolute difference between the mean predictions and the truth.
@@ B.2 Star
Both fits are visually less accurate than the OLAGP and MJGP fits in Figure 9. The global GP performs particularly bad in this case, likely because the proportion of observations in component 1 is much larger than the proportion of values in component 2. The table of RMSPE values in the right panel verifies our intuition, showing that OLAGP and MJGP produce more accurate predictions than either method, particularly the global GP. The LAGP fit in the center panel also shows the selected neighborhoods for the same two predictive locations used in Section 5.2, $(0.15,0.8)$ and $(0.83,0.11)$. Notice in particular that the neighborhood for $(0.15,0.8)$ contains points in both the orange and yellow regions, which is why our prediction is inaccurate. We are able to select smaller and differently shaped neighborhoods using OLAGP and MJGP so that our predictions are more accurate. We can verify that OLAGP and MJGP make more accurate predictions by looking at Figure 18, which shows the absolute difference between the mean predictions and the truth.
@@ C.1 Data with Small Jumps
For the case with small jumps, we modify the function used in Section 1. We define the same jumping sine curve, but this time with a jump of 2 instead of 10. In other words, we use the function in Eq. 1, but replacing the 10 with a 2. We consider modeling this function using three of the four methods in Figure 1; we exclude MJGP with a global GP. The results are shown in Figure 19.
Figure 19: Predictive mean and 95% confidence intervals for the small jump case using a (left) global GP, (center) LAGP, and (right) MJGP.
The black lines represent the true function, while the colors (yellow, red, and blue) represent the predictive mean and 95% prediction intervals. We see that MJGP is still able to accurately capture the jumps, and offers improvement over the global GP and LAGP methods. Calculating RMSPE for each method confirms our findings; the global GP, LAGP, and MJGP produce RMSPEs of 0.8675, 0.4129, and 0.1038, respectively.
@@ C.2 Data with No Jumps
Figure 20: Predictive mean and 95% confidence intervals for the no jump case using a (left) global GP, (center) LAGP, and (right) MJGP.
Here we see that the results do not look much different for any of the three methods. Since there are no jumps in the data, we do not expect using MJGP to offer any improvements. However, it is worth noting that using MJGP did not hurt us; the predictions made by MJGP are just as accurate as those made by LAGP. We confirm this statement using RMSPE; the global GP, LAGP, and MJGP produce RMSPEs of 0.0644, 0.0017, and 0.0015, respectively. This tells us that the method can be used when jumps are suspected but not confirmed without risk of overfitting.
Appendix D Computational Time
@@
Appendix E Prediction Interval Coverage

### v3

@@ 4.2 A new feature for GP regression
This new variable should be related to component membership. We propose using the probability of component membership [Section 4.1], which should be high for observations likely to be in one component and low for the other. In our example, classifier $\hat{f}(x)$ predicts the probability each location $x$ belongs to the red region; it is large (close to 1) in deeply red regions and small (close to 0) in deeply yellow regions. Suppose we treat those values as a new feature, comprising a new column of $X_{N}$ for GP modeling. In the new, higher dimensional input space, points in the red region will be closer together than those from the yellow region, and vice versa. Under the likelihood (2) and when predicting (4), points from the same component will have more influence on each other than on points from opposite components.
@@
To illustrate the effect this new feature has, consider the center plot of Figure 7, again focusing on x=(0.32,0.17)x=(-0.32,0.17) and its nearest neighbor $x_{*}$. We also plot the 160 NNs to $x$, color coded by the rank of their distance from $x$. The points that are the lightest are closest to $x$, and those that are the darkest are furthest away. Note that these distances are calculated prior to hyperparameter fitting. Observe first that the neighborhood is no longer circular; it has become oval-shaped to avoid the yellow region. Any points that are in the yellow region are much further away than their 2d distance would suggest. Their colors are darker than nearby points the red region, and thus farther from $x$.
@@
The right panel of Figure 7 provides a MJGP fit on the Phantom data, contrasting with Figure 4. In this case the GP used is an OLAGP, chosen to illustrate the change in neighborhood structure induced by the MJGP. Here, in Figure 7, both neighborhoods are smaller, but are completely contained to the same side of the manifold. MJGP has improved prediction both visually and via RMSPE (see Appendix B.1) over the global GP, LAGP, and OLAGP.
Now return to Figure 5, summarizing all 10,201 $\hat{n}(x)$ for OLAGP and MJGP. Again, the choice of $n=50$ (the LAGP default) is not particularly advantageous, except for a small percentage of $x\in\mathcal{X}$. Notice that MJGP slightly prefers more mid-range neighborhood sizes than the optimal neighbors GP. In particular, many fewer small $\hat{n}(x)$ are chosen since the neighborhood shape is no longer circular in the original input space (i.e., ignoring the new feature). This facilitates excluding points from the opposite component without resorting to extremely small neighborhood sizes.
@@
We chose to use the predicted probability of component membership $\hat{f}(X_{N})$$\hat{f}(\mathcal{X})$ as a new jump feature. However, we could have instead used the predicted component membership to partition the input space into two regions, and fit separate GPs to each region. If these partitions perfectly matched the manifold of discontinuity, this would outperform MJGP. The problem occurs when the manifold of discontinuity is complex, like in the Phantom example. In this case (see Appendix D), the classifier cannot accurately learn the manifold of discontinuity. MJGP improves upon a partitioned GP by using the predicted probability of component membership, which incorporates the uncertainty of the classifier. Accounting for this extra uncertainty in the model ultimately improves prediction over partitioned GPs.
@@ 5 Implementation and empirical benchmarking
All experiments were completed in R. Code is provided at https://bitbucket.org/gramacylab/jumpgp. In this code we provide functions aGP.optneigh.par(), mjgp(), and jgp() to run OLAGP, MJGP, and OJGP, respectively. OLAGP is parallelized over predictive locations $x\in\mathcal{X}$, turning statistical independence into computational exclusivity via foreach (Microsoft and Weston,, 2022). OJGP leveraged R.matlab (Bengtsson,, 2022).
@@
For MJGP, we utilize EM-based clustering and either a CGP (Williams and Barber,, 1998) or a random forest (RF; Breiman,, 2001) for classification via mclust (Scrucca et al.,, 2023), kernlab (Karatzoglou et al.,, 2004) and randomForest (Liaw and Wiener,, 2002), respectively. We only prefer a RF over a CGP when $n$ is large, for example during the Modified Michalewicz experiment. Again, see Table 1. Model performance is measured out-of-sample using $\mathrm{RMSPE}=(\frac{1}{m}\sum_{x\in\mathcal{X}}(y(x)-\mu(x))^{2})^{1/2},$ for which lower is better. Computation times are in Appendix E and prediction interval coverage is in Appendix F.
@@ 5.1 Phantom
For a more granular view, the right panel draws black lines connecting RMSPEs from the same rep, allowing us to see the best performing method for every training/testing set combination. We conclude that most of the overlap of the boxplots is due to variability in the training/testing split. The ordering of accuracy from one MC rep to the next is largely consistent across methods.
In these plots (and all plots in this section), our two methods are shown third from right (OLAGP) and furthest right (MJGP). Both show improvement over the global GP and LAGP, although OLAGP is typically outperformed by OJGP. This is unsurprising, since OJGP was designed specifically for data from jump processes, and OLAGP was not. But notice that MJGP consistently outperforms OJGP, here and broadly over the next few subsections. Looking first at the line plot on the right of Figure 8, we see that for every data set split, predictions made by MJGP are more accurate than those from OJGP. Additionally, the boxplot shows that OJGP has higher variability than MJGP. OJGP makes predictions by estimating a boundary line separating points from the same and opposite components as the predictive location. Our choice of quadratic boundary yields more accurate predictions, but at the expense of higher variability. For more on this distinction, see Park, (2022).
@@ 5.2 Star
Now consider the “Star” data, which we created for this paper. The goal was to have something like the Phantom, but more angular. Details for how the data set was created are provided in Appendix A.2. See the left panel of Figure 9, with marginal response histogram depicted earlier in the left panel of Figure 6. Like the Phantom, the Star has two well-separated regimes that form a clear shape on the response surface. But the Star is more difficult to model because all points in the yellow regime are geographically close to red ones. It is hard to build a neighborhood inside the yellow portion of the Star.
@@
Repeating themes from the Phantom analysis, consider model performance for a single Star train–test split. Mean fits for the global GP and LAGP are shown in Figure 17 in Appendix B.2, whereas those of OLAGP and MJGP are shown here in the right two panels of Figure 9. RMSPEs can be found in Appendix B.2, where we see that the local methods perform well, with MJGP producing the best predictions of the four. The scenarios depicted in right two panels focus on $x=(0.15,0.8)$ and $(0.83,0.11)$. Compared to LAGP, OLAGP has chosen a neighborhood of roughly the same size for $(0.83,0.11)$, and much smaller for $(0.15,0.8)$; however, this smaller neighborhood still contains points from the red regime. MJGP provides an interesting change: a much smaller neighborhood size for $(0.15,0.11)$, and a larger neighborhood size for $(0.83,0.11)$. Both neighborhoods have adapted their shape to only include points from the same regime, and this improves prediction overall.
@@ 5.3 Modified Michalewicz
Since $N$ here is very large, we must omit the global GP as a competitor in this experiment. Instead of a CGP we deploy a RF as our classifier. Finally, we limit the number of MC reps to thirty. Results are provided in Figure 12. LAGP, OLAGP, and MJGP all considerably outperform OJGP. LAGP and OLAGP perform comparably, although interestingly LAGP seems to slightly outperform OLAGP in this case. We suspect this is due to the spiky nature of the response surface; see Figure 11. In particular, we believe that OLAGP tends to overestimate the optimal neighborhood size for low values of the response (i.e., the spikes), yielding predictions much higher than the true response. MJGP performs best of all, yielding the lowest RMSPE in 29 of the 30 reps.
@@ 5.4 AMHV transport
Finally, consider the AMHV transport data first introduced in Section 1. Specifically, we used a larger version of the data set described in Section 5.2 of Park et al., (2023), scaled so that the inputs lie in the unit hypercube; $d=4$ and N=9,503N=9{,}503. A visual is provided in the right panel of Figure 11. Although $N$ is similar in magnitude to the sample size of the Star and Phantom data sets (N=10,201)(N=10{,}201), the addition of two dimensions means that there is a larger volume to fill. In that sense, this example is the most sparse of the ones entertained in this paper. Outputs are derived from a simulation of the behavior of AMHVs in a particular facility, tallying the average transport time of the AMHVs, including any time spent waiting for the vehicle to arrive. Park et al., (2023) show that as production in the facility increases, transport time, which depends on various design variables, increases and becomes a major bottleneck to the entire manufacturing process. The four variables here include vehicle speed, vehicle acceleration, minimum distance required between vehicles, and the maximum search range of empty vehicles.
@@
The results of the experiment are shown in Figure 13. As expected, the local methods do not perform as well as the global ones, with OLAGP suffering most. But the global GP performs well, producing comparable results to OJGP. MJGP produces only slightly better results, but is consistent in doing so (lowest RMSPEs in 99 of the 100 reps).
@@ B.1 Phantom
Both fits are visually less accurate than the OLAGP fit in Figure 4 and MJGP fit in Figure 7. This is verified by the table of RMSPE values in the right panel, which shows that both OLAGP and MJGP produce more accurate predictions than either method. The LAGP fit in the center panel also shows the selected neighborhoods for the same two predictive locations used in Sections 3 and 4, (0.32,0.17)(-0.32,0.17) and $(0.39,0.34)$. Notice in particular that the neighborhood for $(0.39,0.34)$ contains points in both the red and yellow regions, which explains why our prediction is inaccurate. We are able to select smaller and differently shaped neighborhoods using OLAGP and MJGP so that our predictions are more accurate. We can verify that OLAGP and MJGP make more accurate predictions by looking at Figure 16, which shows the absolute difference between the mean predictions and the truth.
@@ B.2 Star
Both fits are visually less accurate than the OLAGP and MJGP fits in Figure 9. The global GP performs particularly bad in this case, likely because the proportion of observations in component 1 is much larger than the proportion of values in component 2. The table of RMSPE values in the right panel verifies our intuition, showing that OLAGP and MJGP produce more accurate predictions than either method, particularly the global GP. The LAGP fit in the center panel also shows the selected neighborhoods for the same two predictive locations used in Section 5.2, $(0.15,0.8)$ and $(0.83,0.11)$. Notice in particular that the neighborhood for $(0.15,0.8)$ contains points in both the red and yellow regions, which is why our prediction is inaccurate. We are able to select smaller and differently shaped neighborhoods using OLAGP and MJGP so that our predictions are more accurate. We can verify that OLAGP and MJGP make more accurate predictions by looking at Figure 18, which shows the absolute difference between the mean predictions and the truth.
@@ C.1 Data with Small Jumps
For the case with small jumps, we modify the function used in Section 1. We define the same jumping sine curve, but this time with a jump of 2 instead of 10. In other words, we use the function in Eq. 1, but replace the 10 with a 2. We consider modeling this function using three of the four methods in Figure 1; we exclude MJGP with a global GP. The results are shown in Figure 19.
Figure 19: Predictive mean and 95% PI for the small jump case using a (left) global GP, (center) LAGP, and (right) MJGP.
The black lines represent the true function, while the colors (purple, red, and blue) represent the predictive mean and 95% PI. We see that MJGP is still able to accurately capture the jumps, and offers improvement over the global GP and LAGP methods. Calculating RMSPE for each method confirms our findings; the global GP, LAGP, and MJGP produce RMSPEs of 0.1327, 0.4139, and 0.1038, respectively.
@@ C.2 Data with No Jumps
Figure 20: Predictive mean and 95% PI for the no jump case using a (left) global GP, (center) LAGP, and (right) MJGP.
Here we see that the results do not look much different for any of the three methods. Since there are no jumps, we do not expect using MJGP to offer any improvements. However, it is worth noting that using MJGP did not hurt us; the predictions made by MJGP are just as accurate as those made by LAGP. We confirm this statement using RMSPE; the global GP, LAGP, and MJGP produce RMSPEs of 0.0678, 0.0017, and 0.0015, respectively. This tells us that the method can be used when jumps are suspected but not confirmed without risk of overfitting.
Appendix D Comparison to Partitioned GP
One alternative to MJGP is to partition the input space and then fit separate GPs to each partition. However, we prefer MJGP because these partitions are often difficult for a classifier to learn. Consider again the Phantom data set as an example. In Section 4.1 we fit a classifier mapping $X_{N}$ to $C_{N}$, the likely component membership. Rather than using the predicted probability of component membership as a new feature, we could have partitioned the input space into two regions according to their predicted component. The results of such a classification are shown in the left panel of Figure 21.
Figure 21: Left: Predicted component membership by classification model; Center: Global GP fit on partitioned input space; Right: OLAGP fit on partitioned input space.
The manifold of discontinuity is overlaid as a dashed red line. Here we see that the classification does a decent job of matching the outer circle, but struggles to classify the inner circle and ellipses. The top circle is completely misclassified.
Say we were to fit separate GPs based on this partitioning. The results of this method using a global GP and OLAGP are shown in the center and right panels of Figure 21, respectively. In both fits we see the effects of fitting separate GPs; there are jumps along the contours of the classification, even for points on the same side of the manifold of discontinuity. OLAGP offers some improvements over a global GP, but neither performs as well as MJGP; the global GP and OLAGP fits have RMSPEs of 0.1404 and 0.1290, respectively. This is because MJGP is able to incorporate the uncertainty of the classification model, while partitioned GPs ignore this uncertainty.
Appendix E Computational Time
@@
Appendix F Prediction Interval Coverage